OpenStudy (anonymous):

Is the set of numbers that can be written as the product of 6 and an integer closed under subtraction? A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

OpenStudy (anonymous):

@TheEdwardsFamily @Directrix

OpenStudy (anonymous):

help me please

Directrix (directrix):

We need to get an understanding of what these number look like. >Is the set of numbers that can be written as the product of 6 and an integer closed The integers are {...-3, -2, -1. 0, 1, 2, 3, ...}

Directrix (directrix):

6 times each of those would be: {..., -18, -12, -6, 0, 6, 12, 18, ...}

OpenStudy (anonymous):

yes...

Directrix (directrix):

The question is: if you pick any two of those numbers and subtract them, will you get an answer that is also in that set.

OpenStudy (anonymous):

uhm. like..... 6-6=0?

Directrix (directrix):

Yes, because the difference of any two multiples of 6 is equal to 6. 12 is a multiple of 6 24 is a multiple of 6. 12 - 24 is not equal to 6. So, thrown out option A. It is wrong.

OpenStudy (anonymous):

ok

Directrix (directrix):

C. No, and a counterexample is 6 – 6 = 0. 6 - 6 = 0 which is a multiple of 6 because 0 times 6 = 0. So, throw out option C.

OpenStudy (anonymous):

ok

Directrix (directrix):

So, which answer do you think is correct?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

i think....wait. no....D

Directrix (directrix):

Let's see: No, and a counterexample is 6 – 18 = –12. 6 - 18 = -12 but -12 is a multiple of 6 because 6 times -2 = -12. Throw out option D.

Directrix (directrix):

What is left of the options to be the correct one. The option we have not yet investigated.

OpenStudy (anonymous):

B

Directrix (directrix):

B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. Correct Answer.

OpenStudy (anonymous):

OK.... THASNK YOU SOOO MUCH!!!!

Directrix (directrix):

You are welcome.